OpenStudy (anonymous):

What would the Graph for these equations look like? y − 2x = −1 x + 3y = 4 Please help :)

OpenStudy (anonymous):

Ok. I'll post the graph in a few minutes :D

OpenStudy (gorv):

y=2x-1 at x=0 y=-1 (0,-1) at y=0 , x=1/2 (1/2,0)

OpenStudy (gorv):



OpenStudy (anonymous):

Well, here is the first graph. I give you the equations I used. You first need to change into y=mx+b form. Here are the two equations y=-1/3x+4/3 and y=2x-1. Hopes it helps:D

OpenStudy (gorv):

x+3y=4 3y=-x+4 y=-x/3 +4/3 at x=0 y =4/3 (0,4/3) at y=0 x=4 (4,0)

OpenStudy (gorv):